# Generalized Heisenberg Dynamics Revisited

### v1

@@ Appendix A Details of derivation
××$~{}~{}~{}~{}\times\cdots\times((E_{a_{n-1}})_{l_{n-1}~{}\!\!l_{n+1}}-(E_{a_{n-1}})_{l_{n}~{}\!\!l_{n+1}})$
@@
×$~{}~{}~{}~{}\times\prod_{r=1}^{k-1}(E_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(E_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}},$ (157)
where the factor $(-1)^{(n-k)(k-1)}$ appears from the $(n-k)(k-1)$ exchange of the indices on $E_{a_{r}}$ from $(a_{1},a_{2},\cdots,a_{n-k},a_{n-k+1},a_{n-k+2},\cdots,a_{n-1})$ to $(a_{k},a_{k+1},\cdots,a_{n-1},a_{1},a_{2},\cdots,a_{k-1})$. By inserting eq. (157) into eq. (112), eq. (113) is obtained as follows,
@@
where we use $(n-1)k+(n-k)(k-1)=2nk-k(k+1)-(n-k)=n-k$ (mod 2). From the fact that eq. (109) holds for $(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n}}$, it is shown that $(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n}}$ satisfies the rule (114) as follows,
@@
×$~{}~{}~{}~{}\times((E_{a_{n-p-1}})_{l_{n-1}~{}\!\!l_{n}}+(E_{a_{n-p-1}})_{l_{n}~{}\!\!l_{p}})(E_{a_{n-p}})_{l_{n}~{}\!\!l_{p}}$
×$~{}~{}~{}~{}\times((E_{a_{n-p+1}})_{l_{1}~{}\!\!l_{n}}+(E_{a_{n-p+1}})_{l_{n}~{}\!\!l_{p}})\cdots((E_{a_{n-1}})_{l_{p-1}~{}\!\!l_{n}}+(E_{a_{n-1}})_{l_{n}~{}\!\!l_{p}})$
@@
where the factor $-(-1)^{(n-1)p}$ comes from the $(n-2)(n-p-1)+(p-1)(=(n-1)p+1({\rm mod}~{}2))$ exchange of the indices on $E_{a_{r}}$ from $(a_{n-p+1},\cdots,a_{n-1},a_{n-p},a_{1},a_{2},\cdots,a_{n-p-1})$ to $(a_{1},\cdots,a_{n-p-2},a_{n-p-1},a_{n-p},a_{n-p+1},\cdots,a_{n-1})$.
@@
×$~{}~{}~{}~{}~{}\times\prod_{r=1}^{k-1}(h_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}}$ (162)
@@
××$~{}~{}~{}~{}~{}~{}\times\cdots\times(H_{2})_{l_{1}~{}\!\!l_{2}~{}\!\!l_{n+1}~{}\!\!l_{4}~{}\!\!\cdots~{}\!\!l_{n}}(H_{1})_{l_{1}~{}\!\!l_{n+1}~{}\!\!l_{3}~{}\!\!\cdots~{}\!\!l_{n}}(H_{n-1})_{l_{n+1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$
@@
××$~{}~{}~{}~{}~{}~{}\times\cdots\times(H_{2})_{l_{1}~{}\!\!l_{2}~{}\!\!l_{n}~{}\!\!l_{4}~{}\!\!\cdots~{}\!\!l_{n}}(H_{1})_{l_{1}~{}\!\!l_{n}~{}\!\!l_{3}~{}\!\!\cdots~{}\!\!l_{n}}(H_{n-1})_{l_{n}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$
@@
×$~{}~{}~{}~{}~{}~{}~{}~{}\times A(t)_{l_{1}~{}\!\!\cdots~{}\!\!l_{k-1}~{}\!\!l_{n+1}~{}\!\!l_{k+1}~{}\!\!\cdots~{}\!\!l_{n}}(H_{k-2})_{l_{1}~{}\!\!\cdots~{}\!\!l_{k-2}~{}\!\!l_{n+1}~{}\!\!l_{k}~{}\!\!\cdots~{}\!\!l_{n}}\cdots(H_{2})_{l_{1}~{}\!\!l_{2}~{}\!\!l_{n+1}~{}\!\!l_{4}~{}\!\!\cdots~{}\!\!l_{n}}(H_{1})_{l_{1}~{}\!\!l_{n+1}~{}\!\!l_{3}~{}\!\!\cdots~{}\!\!l_{n}}$
×$~{}~{}~{}~{}~{}~{}~{}~{}\times(H_{n-1})_{l_{n+1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$
@@
×$~{}~{}~{}~{}~{}~{}~{}~{}\times(H_{k-2})_{l_{1}~{}\!\!\cdots~{}\!\!l_{k-2}~{}\!\!l_{k}~{}\!\!l_{k}~{}\!\!\cdots~{}\!\!l_{n}}\cdots(H_{2})_{l_{1}~{}\!\!l_{2}~{}\!\!l_{k}~{}\!\!l_{4}~{}\!\!\cdots~{}\!\!l_{n}}(H_{1})_{l_{1}~{}\!\!l_{k}~{}\!\!l_{3}~{}\!\!\cdots~{}\!\!l_{n}}(H_{n-1})_{l_{k}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$
@@
for $k\leq j\leq n-1$. In eqs. (174), (175), (176) and (177), the $n-1$ indices $(l_{1},\cdots,l_{j},l_{j+2},\cdots l_{n})$, $(l_{1},\cdots,l_{k-2},l_{k},\cdots l_{n})$, $(l_{1},\cdots,l_{k},l_{k+2},\cdots l_{n})$, and $(l_{1},\cdots,l_{j+1},l_{j+3},\cdots l_{n})$ are all different, respectively. The ellipsis in eq. (173) stands for terms including both $(h_{a_{k}})_{l_{r}~{}\!\!l_{n}}$ and $(h_{a_{k^{\prime}}})_{l_{r}~{}\!\!l_{n}}$ ($a_{k}\neq a_{k^{\prime}}$) and all of them are eliminated by adding the terms which stem from the exchange of $H_{a_{k}}$ and $H_{a_{k^{\prime}}}$.
@@
×$~{}~{}~{}\times\prod_{r=1}^{k-1}(h_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}}~{}\!A(t)_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}},$ (182)

### v2

@@ Appendix A Details of derivation
$~~~~\times\cdots\times((E_{a_{n-1}})_{l_{n-1}~\!\!l_{n+1}}-(E_{a_{n-1}})_{l_{n}~\!\!l_{n+1}})$
@@
$~~~~\times\prod_{r=1}^{k-1}(E_{a_{r}})_{l_{r}~\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(E_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~\!\!l_{k}},$ (168)
where the factor $(-1)^{(n-k)(k-1)}$ appears from the (nk)(k1)(n-k)(k-1) exchange of the indices on $E_{a_{r}}$ from $(a_{1},a_{2},\cdots,a_{n-k},a_{n-k+1},a_{n-k+2},\cdots,a_{n-1})$ to $(a_{k},a_{k+1},\cdots,a_{n-1},a_{1},a_{2},\cdots,a_{k-1})$. By inserting eq. (168) into eq. (123), eq. (124) is obtained as follows,
@@
where we use (n1)k+(nk)(k1)=2nkk(k+1)(nk)=nk(n-1)k+(n-k)(k-1)=2nk-k(k+1)-(n-k)=n-k (mod 2). From the fact that eq. (120) holds for $(\nu_{n})_{l_{1}~\!\!l_{2}~\!\!\cdots l_{n}}$, it is shown that $(\nu_{n})_{l_{1}~\!\!l_{2}~\!\!\cdots l_{n}}$ satisfies the rule (125) as follows,
@@
$~~~~\times((E_{a_{n-p-1}})_{l_{n-1}~\!\!l_{n}}+(E_{a_{n-p-1}})_{l_{n}~\!\!l_{p}})(E_{a_{n-p}})_{l_{n}~\!\!l_{p}}$
$~~~~\times((E_{a_{n-p+1}})_{l_{1}~\!\!l_{n}}+(E_{a_{n-p+1}})_{l_{n}~\!\!l_{p}})\cdots((E_{a_{n-1}})_{l_{p-1}~\!\!l_{n}}+(E_{a_{n-1}})_{l_{n}~\!\!l_{p}})$
@@
where the factor $(-1)^{(n-1)p}$ comes from the $(n-2)(n-p-1)+(p-1)(=(n-1)p+1({\rm mod}~2))$ exchange of the indices on $E_{a_{r}}$ from $(a_{n-p+1},\cdots,a_{n-1},a_{n-p},a_{1},a_{2},\cdots,a_{n-p-1})$ to $(a_{1},\cdots,a_{n-p-2},a_{n-p-1},a_{n-p},a_{n-p+1},\cdots,a_{n-1})$.
@@
$~~~~~\times\prod_{r=1}^{k-1}(h_{a_{r}})_{l_{r}~\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~\!\!l_{k}}$ (173)
@@
$~~~~~~\times\cdots\times(H_{2})_{l_{1}~\!\!l_{2}~\!\!l_{n+1}~\!\!l_{4}~\!\!\cdots~\!\!l_{n}}(H_{1})_{l_{1}~\!\!l_{n+1}~\!\!l_{3}~\!\!\cdots~\!\!l_{n}}(H_{n-1})_{l_{n+1}~\!\!l_{2}~\!\!\cdots~\!\!l_{n}}$
@@
$~~~~~~\times\cdots\times(H_{2})_{l_{1}~\!\!l_{2}~\!\!l_{n}~\!\!l_{4}~\!\!\cdots~\!\!l_{n}}(H_{1})_{l_{1}~\!\!l_{n}~\!\!l_{3}~\!\!\cdots~\!\!l_{n}}(H_{n-1})_{l_{n}~\!\!l_{2}~\!\!\cdots~\!\!l_{n}}$
@@
$~~~~~~~~\times A(t)_{l_{1}~\!\!\cdots~\!\!l_{k-1}~\!\!l_{n+1}~\!\!l_{k+1}~\!\!\cdots~\!\!l_{n}}(H_{k-2})_{l_{1}~\!\!\cdots~\!\!l_{k-2}~\!\!l_{n+1}~\!\!l_{k}~\!\!\cdots~\!\!l_{n}}\cdots(H_{2})_{l_{1}~\!\!l_{2}~\!\!l_{n+1}~\!\!l_{4}~\!\!\cdots~\!\!l_{n}}(H_{1})_{l_{1}~\!\!l_{n+1}~\!\!l_{3}~\!\!\cdots~\!\!l_{n}}$
$~~~~~~~~\times(H_{n-1})_{l_{n+1}~\!\!l_{2}~\!\!\cdots~\!\!l_{n}}$
@@
$~~~~~~~~\times(H_{k-2})_{l_{1}~\!\!\cdots~\!\!l_{k-2}~\!\!l_{k}~\!\!l_{k}~\!\!\cdots~\!\!l_{n}}\cdots(H_{2})_{l_{1}~\!\!l_{2}~\!\!l_{k}~\!\!l_{4}~\!\!\cdots~\!\!l_{n}}(H_{1})_{l_{1}~\!\!l_{k}~\!\!l_{3}~\!\!\cdots~\!\!l_{n}}(H_{n-1})_{l_{k}~\!\!l_{2}~\!\!\cdots~\!\!l_{n}}$
@@
for $k\leq j\leq n-1$. In eqs. (185), (186), (187) and (188), the n1n-1 indices $(l_{1},\cdots,l_{j},l_{j+2},\cdots l_{n})$, $(l_{1},\cdots,l_{k-2},l_{k},\cdots l_{n})$, $(l_{1},\cdots,l_{k},l_{k+2},\cdots l_{n})$, and $(l_{1},\cdots,l_{j+1},l_{j+3},\cdots l_{n})$ are all different, respectively. The ellipsis in eq. (184) stands for terms including both $(h_{a_{k}})_{l_{r}~\!\!l_{n}}$ and $(h_{a_{k^{\prime}}})_{l_{r}~\!\!l_{n}}$ ($a_{k}\neq a_{k^{\prime}}$) and all of them are eliminated by adding the terms which stem from the exchange of $H_{a_{k}}$ and $H_{a_{k^{\prime}}}$.
@@
$~~~\times\prod_{r=1}^{k-1}(h_{a_{r}})_{l_{r}~\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~\!\!l_{k}}~\!A(t)_{l_{1}~\!\!l_{2}~\!\!\cdots~\!\!l_{n}},$ (193)
@@
Appendix B Reduction to the Heisenberg dynamics
In the subsection 3.1, we have seen how Nambu dynamics is reduced to Hamiltonian dynamics. In concrete, the Nambu equation turns out to be Hamilton’s canonical equations of motion by reducing the phase space.
In this appendix, we show that the generalized Heisenberg equation can be reduced to Heisenberg’s equation of motion by taking a specific form of Hamiltonians. The starting point is the extension of Heisenberg’s equation of motion:
$\frac{d}{dt}A^{\alpha}_{l~\!\!m~\!\!n}(t)=\frac{1}{i\hbar}\left[A^{\alpha}(t),H_{1},H_{2}\right]_{l~\!\!m~\!\!n},$ (196)
where $A^{\alpha}_{l~\!\!m~\!\!n}(t)$ ($l,m,n=0,1,2\cdots,N$, $\alpha=1,2,3$) are the components of $(N+1)\times(N+1)\times(N+1)$ generalized matrices $A^{\alpha}=(X,Y,Z)$, $H_{1}$ and $H_{2}$ are Hamiltonians given by $(N+1)\times(N+1)\times(N+1)$ generalized matrices. We note that $N$ can be infinity depending on a system.
In general, $H_{1}$ depends on some variables of $A^{\alpha}$ and it takes specific values depending on the solution. Let us suppose that the following solution is given by
$X_{l~\!\!m~\!\!n}(t)=\begin{cases}X_{l~\!\!m~\!\!n}(0)e^{2\pi i\nu_{m~\!\!n}t}&(l=0,m,n\neq 0)\\ 0&({\rm otherwise})\end{cases},$ (197)
$Y_{l~\!\!m~\!\!n}(t)=\begin{cases}Y_{l~\!\!m~\!\!n}(0)e^{2\pi i\nu_{m~\!\!n}t}&(l=0,m,n\neq 0)\\ 0&({\rm otherwise})\end{cases}$ (198)
and $H_{1}$ takes the values such as
$(H_{1})_{l~\!\!m~\!\!n}=\begin{cases}E_{m}\delta_{m~\!\!n}&(l=0,m,n\neq 0)\\ 0&({\rm otherwise}),\end{cases}$ (199)
where $X_{l~\!\!n~\!\!m}(0)=\overline{X_{l~\!\!m~\!\!n}(0)}$, $Y_{l~\!\!n~\!\!m}(0)=\overline{Y_{l~\!\!m~\!\!n}(0)}$ and $h\nu_{m~\!\!n}=E_{m}-E_{n}$.
When $(H_{2})_{l~\!\!m~\!\!n}=1$ for $l,m,n\neq 0$, the equation (196) can be rewritten in the form of Heisenberg’s equation of motion:
$\frac{d}{dt}A^{\alpha}_{l~\!\!m~\!\!n}(t)=\frac{1}{i\hbar}\left[A^{\alpha}(t),H\right]^{\prime}_{l~\!\!m~\!\!n},$ (200)
where $H=H_{1}$ and $\left[A^{\alpha}(t),H\right]^{\prime}_{l~\!\!m~\!\!n}$ is defined by
$\left[A^{\alpha}(t),H\right]^{\prime}_{l~\!\!m~\!\!n}\equiv\sum_{k=1}^{N}\left(A^{\alpha}(t)_{l~\!\!m~\!\!k}H_{l~\!\!k~\!\!n}-H_{l~\!\!m~\!\!k}A^{\alpha}(t)_{l~\!\!k~\!\!n}\right).$ (201)
Furthermore, the commutation relation $X,Y,Z]_{l~\!\!m~\!\!n}=i\hbar I_{l~\!\!m~\!\!n}$ can be reduced to the usual one $X,Y,Z]_{0~\!\!m~\!\!n}=[X,Y]^{\prime}_{0~\!\!m~\!\!n}=i\hbar\delta_{m~\!n}$ for $m,n\neq 0$, if we choose $Z=H_{2}$.
In this way, we find that Heisenberg’s matrix mechanics can be described by the generalized matrices.